OpenStudy (anonymous):

Find the equation of the line that contains the given points. Let y be the dependent variable and let x be the independent variable. P1(3, -9), P2(5, -9) this is what i did and its coming out wrong... y-(-9)=x-3 y+9=x-3 y=x-(3-9) y=x+6 what am I dont wrong?

OpenStudy (saifoo.khan):

you forgot the slope.

OpenStudy (saifoo.khan):

and the formula is, \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

I have the formula down I just added in the numbers

OpenStudy (saifoo.khan):

DId you found the slope?

OpenStudy (anonymous):

m=0

OpenStudy (saifoo.khan):

Then it will be something like, y = -9

OpenStudy (saifoo.khan):

Try graphing it.

OpenStudy (lgbasallote):

then you shouldnt have an x if your m is zero right? coz it's y-y1 = m(x-x1)...if your m is zero...your whole x becomes zero too since you multiply it by zero

OpenStudy (anonymous):

yeah I just saw that

OpenStudy (anonymous):

it works y=-9

OpenStudy (anonymous):

slope is rise over run so slope is how much y changes over how much x changes

OpenStudy (anonymous):

thank you